HUBBLE SPACE TELESCOPE DAILY REPORT #2509 PERIOD COVERED: 0000Z (UTC) 11/29/99 - 0000Z (UTC) 11/30/99 Daily Status Report as of 334/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: Since HST remains in Zero Gyro Sun Point (ZGSP), no science observations were performed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No activity 2.2 FHST Updates: No activity 2.3 Operations Notes: The HST remains in ZGSP. The Goldstone Deep Space Network site was used to dump SSR engineering data to the ground. Operations Note #661 was utilized to reset the time-based recharge ratio charge/discharge buffers at 333/1356Z. The CCS string "G" FEP required a reboot. This was followed by a re-cycle of the "G" string at 333/1843Z. The CCS "B" string was re-cycled at 333/2027Z, and configured for TDRS back-up. Following each re-cycle, numerous limit changes were made. 3.0 SIGNIFICANT FORTHCOMING EVENTS: The PSEA/RMGA On-Orbit Test will be conducted today from 334/1200Z until approximately 334/2100Z. CCS string "A" will be utilized with CCS string "B" configured as hot back-up. Servicing Mission 3A, currently scheduled for December 9, 1999. /CAW